OpenStudy (anonymous):

g(x) = x^2 - 4x- 12 What are the x intercepts?

OpenStudy (anonymous):

(x-2)^2)-16 (x-6)(x+2)

OpenStudy (anonymous):

at first i added -4 and +4 for this question x^2 - 4x- 12 +4 -4 (x^2 - 4x+4)-16 (x-2)^2)-16 (x-6)(x+2)

OpenStudy (amistre64):

the xints occur when g(x) = 0

OpenStudy (anonymous):

x=-2 x=6 proof x^2-4x-12=0 x-intercepts means y=0 comparing with general equation ax^2+bx+c=0 12*1=12 factors of 12 used = 2 , 6 hence, x^2-6x+2x-12=0 x(x-6)+2(x-6)=0 (x-6)(x+2)=0 therefore x=-2 x=6

OpenStudy (anonymous):

y=2/3x-3.............eqn 1 2x-3y=9............eqn 2 put eqn 1 into eqn 2 2x-3(2/3x-3)=9 2x-2x+9=9 2x-2x+9-9=0 hence x and y has no values